Write the following statements in the form of equation or expression , you are not required to solve the equation:
a). 4 more than 5 times a number , plus 2 times that number increased by 40
b). 21 less than twice c gives 9
c). 5 more than 6 times a number, plus 3 times that number, is 25.
d). A number less than 24 is the average of 15, 5 , 32 ?

Answer by mananth(16946)

Write the following statements in the form of equation or expression , you are not required to solve the equation:

a). 4 more than 5 times a number , plus 2 times that number increased by 40
5x+4 +2x+40
b). 21 less than twice c gives 9
2c-21=9
c). 5 more than 6 times a number, plus 3 times that number, is 25.
6x+5 + 3x=25
d). A number less than 24 is the average of 15, 5 , 32 ?
24-x = (15+5+32)/3